what is the value of z? -3z+9(z+2)-4=18
Answer by mananth(16946)

-3z+9(z+2)-4=18
-3z+9z+18-4=18
6z+14=18
-14
6z+14-14=18-14
6z=4
/6
z=4/6 =2/3